OpenStudy (anonymous):

y = 2 - x -5x - 15y = 60 Which method would you use to solve this system? What is the solution to this system of equations?

OpenStudy (anonymous):

get all variables(x and y) on one side and the constants to the other side(2 and 60).. then multiply one of them so as to have one of the coefficients same (for eg, you can multiply equation one with 5, so as to have the same coefficients for x ) then subtract or add, in order to eliminate that variable ( in our eg, eliminate x), and solve to get the answer for the other variable, finally substitute and get value of the eliminated variable (x)

OpenStudy (anonymous):

multiply first equation by 5 and add the two :)

OpenStudy (anonymous):

Substitution Substitute y = (2 - x) in second equation -5x - 15(2 - x)=60 -5x-30+15x=60 10x=90 x=9 put x=9 in y = 2 - x and simplify y=2-9=-7 solution x=9, y=-7